### v1

# Defending against adversarial attacks using mixture of experts

Abstract
Machine learning is a powerful tool enabling full automation of a huge number of tasks without explicit programming. Despite recent progress of machine learning in different domains, these models have shown vulnerabilities when they are exposed to adversarial threats. Adversarial threats aim to hinder the machine learning models from satisfying their objectives. They can create adversarial perturbations, which are imperceptible to humans’ eyes but have the ability to cause misclassification during inference. Moreover, they can poison the training data to harm the model’s performance or they can query the model to steal its sensitive information. In this paper, we propose a defense system, which devises an adversarial training module within mixture-of-experts architecture to enhance its robustness against adversarial threats. In our proposed defense system, we use nine pre-trained experts with ResNet-18 as their backbone. During end-to-end training, the parameters of expert models and gating mechanism are jointly updated allowing further optimization of the experts. Our proposed defense system outperforms state-of-the-art defense systems and plain classifiers, which use a more complex architecture than our model’s backbone.
keywords:
Adversarial Machine Learning, Mixture of Experts, Robustness, Ensemble Learning, Adversarial Training, and Evasion Attacks.
journal: Information Sciences
[…]
organization=Computer Science, University of New Brunswick, city=Fredericton, postcode=E3B 5A3, country=Canada
1 Introduction
Machine learning (ML) is a cornerstone of artificial intelligence (AI), which has been widely used in numerous applications. In computer vision, it enables face recognition and video surveillance [khan2021machine, khan2020machine]. In health care, it facilitates disease diagnosis, drug repurposing, and personalized treatments [habehh2021machine, qayyum2020secure]. In natural language processing (NLP), ML enhances the quality of language translation [lauriola2022introduction, omar2022robust]. ML also empowers Large Language Models (LLMs) for code-related tasks such as code generation and understanding [jelodar2025large]. In cybersecurity, ML eases threat detection and manages system’s vulnerabilities [apruzzese2023role, xin2018machine, buczak2015survey]. Moreover, ML also plays an important role in other domains such as finance, speech processing, recommendation systems, autonomous vehicles, among others. However, those wide range of applications of machine learning has introduced opportunities for adversaries to exploit natural vulnerabilities of these models and launch adversarial attacks to threaten the integrity, privacy, and availability of these models [pitropakis2019taxonomy, bountakas2023defense, kurakin2016adversarial, kaur2022trustworthy].
Adversarial machine learning (AML) is at intersection of cybersecurity and machine learning and aims to investigate attacks on machine learning models, and design defense systems to counter these attacks. The adversary may have some knowledge about the victim’s model and/or its training data, which facilitates the adversarial sample generation [pitropakis2019taxonomy, bountakas2023defense, kurakin2016adversarial]. Figure 1 shows how an adversarial sample looks like on an MNIST [lecun2002gradient] sample when the adversary has a perfect knowledge about the model’s parameters. Samples like this can cause model’s misclassification with a high confidence during test time.
Figure 1: Generating an adversarial example on an MNIST digit by adding a carefully crafted perturbation, which remains imperceptible to humans but successfully fools the target models.
The rivalry between governments and industrial corporations has given rise to adversarial attacks causing global concerns about adversarial machine learning. For instance, the government of United States of America has established Bipartisan AI task force, which are responsible for appropriate use of AI. They introduced AI guidelines such as accuracy, reliability, robustness, privacy, transparency, and safety. They also labeled generative and adversarial AI as a threat to privacy and confidentiality [AI_Task_Force_2024]. Moreover, the Department of Homeland Security (DHS) has considered evasion attacks and generative AI as a national threat violating security and privacy of individuals [DHS_AI_Threat_2023]. These insights highlight the need for secure AI systems and robust ML models.
Trustworthy AI is an important domain of AI ensuring that AI models remain lawful, ethical, robust, and efficient. Robust machine learning is an important branch in trustworthy AI which aims to make the machine learning models robust against adversarial threats while remain functional in non-adversarial environments [kaur2022trustworthy]. Moreover, the main focus of Robust ML is to develop defense systems that detect and/or mitigate adversarial threats while preserving their efficiency and performance to a certain extent. These defense systems could either reside in the models or serve as an external entity protecting ML models [pitropakis2019taxonomy, bountakas2023defense, kurakin2016adversarial, kaur2022trustworthy].
In pursuit of building a secure system, we build a novel defense system named Divided We Fall (DWF). In DWF, we employ Mixture of Experts (MoE) architecture which consists of benign and adversarial experts. The training scheme in DWF is joint, meaning that we do not freeze the weights of our pretrained experts during end-to-end training, which leads to stronger generalization capability of our proposed defense system improving the clean accuracy and robustness of our model.
Our contributions are as follows:
  • Our proposed defense mechanism devises an adversarial training module within MoE, which outperforms related state-of-the-art defense systems in terms of both clean accuracy and robustness.
  • Our model maintains high accuracy, while significantly outperforms undefended models in terms of robustness.
  • We introduce a new training approach for MoE, where pretrained experts are jointly updated along with the gating mechanism during an end-to-end training session, where we use a mixed batch encompassing both benign and adversarial data.
  • The randomized attack settings during all training stages enable DWF to generalize more efficiently in comparison to its competitors.
In Section 2, we illustrate the background problem setting. In Section 3, we introduce similar defense systems. In Section 4, we present the methodology and workflow of our defense. In Section 5, we demonstrate our experimental results. In Section 6, we investigate our proposed defense system limitations, potential future directions, and finally conclude our work.
2 Background and Problem Setting
MoE is an ensemble architecture using multiple experts, a gating mechanism to assign weights, and a router to aggregate results [mu2025comprehensive, cai2025survey, puigcerver2023sparse]. MoEs can be categorized based on their gating mechanism, routing topology, training strategy, and experts activation, which is shown in Figure 2.
Based on gating mechanism, MoE could be either soft or hard; in soft MoE the gating assigns non-zero weights to each expert forming fractional activation of experts, while in hard MoE, the gating mechanism chooses a subset of experts and assign wights to them making experts activation as sparse, or single if only one of them is chosen [mu2025comprehensive, cai2025survey, puigcerver2023sparse].
Training strategy could be either joint or frozen. In the first one, both router and experts are trained simultaneously, while the latter one uses pretrained experts and freezes their weights during training of the router [mu2025comprehensive, cai2025survey, puigcerver2023sparse].
Based on the routing topology we have flat and hierarchical routing schemes. In the flat MoE, all the experts reside in a single layer and the router chooses from them. In the hierarchical topology, the first router groups the experts into subgroups, and, then, the second one chooses some experts from these subgroups [mu2025comprehensive, cai2025survey].
Figure 2: Main components and types of mixture of experts.
Adversarial threats can be viewed from different angles: attacker’s knowledge, type, specificity, and objective. Figure 3 shows an overview of adversarial threat model.
2.1 Attacker’s Knowledge
Attacker’s knowledge shows the access level of the adversary to model’s parameters, gradients, architecture, and training data. This knowledge is categorized into white-box, black-box, and gray-box based on the adversary’s access level [pitropakis2019taxonomy, bountakas2023defense].
2.1.1 White Box
Adversaries with white-box knowledge are the most effective and destructive attackers. These attackers have a full access to model’s architecture, parameters, and gradients during training with full knowledge on the training data; this enables the attackers to generate adversarial samples based on the gradients and training data, which makes them very powerful. However, in a real-world scenario, adversaries can’t achieve such an access, making them more suitable for research and testing the models’ robustness [pitropakis2019taxonomy, bountakas2023defense].
2.1.2 Black Box
In black-box scenario, unlike white-box, the adversaries have no knowledge about the model and the presence of any defense system. The adversaries can only give inputs and receive the corresponding outputs [pitropakis2019taxonomy, bountakas2023defense, lin2023sensitive].
2.1.3 Gray Box
Gray box knowledge lies between white box knowledge and black box knowledge. The adversary has some degree of knowledge about the target system, which is not as perfect as white box, and is not as scarce as black box knowledge [pitropakis2019taxonomy, bountakas2023defense].
2.2 Attacker’s specificity
Based on specificity, each attack could either be targeted or untargeted. In targeted attacks, the adversary aims to degrade the model’s performance when it faces specific patterns/inputs, while in untargeted attacks, also named to as indiscriminate, the goal is to cause any misclassification [pitropakis2019taxonomy, bountakas2023defense].
2.3 Attacker’s objective
Based on the objective, the adversary can follow degrade, breach, or both scenarios. In the first scenario, the adversary aims to harm the system’s performance and/or violets its integrity. In breach, the adversary aims to explore the model in order to steal sensitive information, without affecting the model’s performance [pitropakis2019taxonomy, bountakas2023defense].
2.4 Attacker’s type
In a general perspective, an adversarial attack belongs to one of the following types: evasion, poisoning, backdoor, or physical layer attacks.
2.4.1 Evasion
An evasion attack targets the model at test time without affecting the system’s integrity on clean data. Evasion attacks are the most practical attack types aiming to degrade the model’s performance by causing misclassification during testing [pitropakis2019taxonomy, bountakas2023defense, gong2024improving]. For instance attacks such as Fast Gradient Sign Method (FGSM) [goodfellow2014explaining], Projected Gradient Descent (PGD) [madry2017towards], and Carlini & Wagner (C&W) [carlini2017towards] make use of the model’s gradient during training to generate samples which evade the classifier during the test time. Equations 2 and 3, in Section 5, show how an adversarial sample is generated in FGSM and PGD, respectively.
2.4.2 Poisoning
Poisoning attacks aim to violate the system’s integrity by corrupting the training data and/or model embeddings to harm the model integrity. These attacks could be either label poisoning attacks or clean label attacks, where only the former manipulates the labels. Additionally, backdoor poisoning attacks insert hidden triggers/patterns in the training data forcing the model to cause misclassification when some target patterns are observed by the model without affecting model’s overall performance prominently [pitropakis2019taxonomy, bountakas2023defense, tian2022comprehensive, bai2024backdoor, zhang2025apbam, zheng2022concealed].
2.4.3 Exploratory
Exploratory attacks aim to violate privacy of the system. These attacks probe the target model to gain information about model’s parameters, training data, and so on without affecting the performance and integrity of the model [rigaki2023survey, zhang2024defending]. One of the most well-known attacks from this category is named Membership Inference Attack (MIA) [shokri2017membership], which aims to assess the information leakage of a machine learning model by giving an input and receiving the corresponding output in a black-box manner.
2.4.4 Physical
These attacks are launched on the models, which are deployed on a special physical hardware such as FPGA, GPU, edge devices and so on. Attacks on physical layer exploit vulnerabilities, which are emerged due to the increased attack surface in the physical layer [regazzoni2020machine]. For instance, in a power analysis attack, the adversary can infer cryptographic secrets from the amount of power used by a device, since power consumption depends on the amount of data and the number of operations [hettwer2020applications].
Figure 3: General taxonomy of the threat model.
All together, what we mentioned illustrates the seriousness of adversarial threats, highlighting the urge to build robust models to defend these adversarial threats, emphasizing the importance of adversarial machine learning and robust models.
3 Related Works
Given the arms race nature of adversarial machine learning, a plethora of attacks and defenses have been proposed and evolved over the years. This highlights the evolving nature of AML. In the following paragraphs, we investigate the related defenses in this area.
Adversarial training is a defense strategy aiming to train the model on both benign and adversarial samples to increase generalization capability of the model enabling it to correctly predict both benign and adversarial inputs. However, this method often comes with a trade-off, where the model sacrifices it accuracy on the benign data to increase its robustness against adversarial inputs [bai2021recent, zhao2024adversarial, zhang2025catil, jiang2024prior].
Adversarial training was first introduced in [goodfellow2014explaining], which increased the model’s robustness against FGSM attacks by sacrificing the clean accuracy of the model. Methods such as TRadeoff-inspired Adversarial DEfense via Surrogate-loss minimization (TRADES) [zhang2019theoretically] and Fair Robust Learning (FRL) [xu2021robust] have been introduced to control natural robustness-accuracy trade-off in adversarial training. Additionally, methods such as Fast Free Adversarial Training (FFAT) [shafahi2019adversarial] tackles the computational overhead of adversarial sample generation, while having a lower accuracy and robustness in comparison to standard adversarial training.
Ensemble architectures have shown strong robustness against adversarial threats in comparison to the non-ensemble architectures. This is due to the natural complex architecture of ensemble models. Ensemble-based defense methods are categorized into distillation, hierarchical, and MoE.
Ensemble distillation-based defenses aim to harden the model by training it on soft labels and/or noisy logits instead of hard labeling, producing results via ensembles. Research works in this category include Distillation Hardening for Random Forest (DDH-RF) [apruzzese2020hardening] and SELENA [tang2022mitigating]. Hierarchical defenses, on the other hand, aim to build ensemble of models in a hierarchical manner to detect and mitigate adversarial attacks in multiple levels, which make the defense systems robust against a wider range of attacks. Hierarchical defenses include Application Constraints (AppCon) [apruzzese2020appcon], Model Stacking [salem2018ml], and Ensemble Methods as a Defense (EMD) [strauss2017ensemble].
[puigcerver2022adversarial, pavlitska2024towards] investigate the adversarial robustness of MoEs. In [han2024enhancing, zhang2025optimizing], some defense systems are proposed, which aim to protect the architecture of MoEs. Defenses such as adversarial mixture-of-expert (ADVMoE) [zhang2023robust] and synergy-of-experts (SoE) [cui2022synergy] make use of the ensemble architecture of MoEs to defend against adversarial threats. Both ADVMoE and SoE use adversarial training techniques, while training their MoE to increase the robustness of their model. Table 1 presents the differences between the category of MoE used in our defense system and other competitors.
Table 1: MoE type of DWF vs. other competitors - gating mechanism (GM), routing topology (RT), training strategy (TS), and experts activation (EA).
Defense GM RT TS EA
SoE Hard Flat Joint Single
ADVMoE Hard Flat Joint Sparse
DWF Soft Flat Pre-trained+Joint Fractional
4 Methodology
Figure 4: General view of the end to end training of our proposed defense system.
Our defense uses soft and flat structure for MoE with fractional activation of experts. In the first phase of training, we independently train a benign expert, four FGSM experts, and four PGD experts. After this pretraining stage, we move to end-to-end learning, in which the entire framework is jointly optimized. During this stage, for each batch of data, we generate adversarial variants of the inputs using both FGSM and PGD attacks, and augment them with the original benign inputs to form an expanded batch. This mixed batch preserves the diversity of benign and adversarial samples, enhancing the network’s robustness and generalization capability. Finally, the combined batch is fed into the network, so the whole framework learns both benign and adversarial patterns. Figure 4 shows an overview of our defense model’s workflow.
To construct FGSM experts, we need to pre-train them before end-to-end training. During adversarial training of an FGSM expert, within every batch and each epoch we randomly assign a value between 0.01 to 0.09 to perturbation size $\epsilon$, and the models is adversarially trained on random perturbation size during that epoch. This strategy helps to increase generalization ability of the network, making it more robust against different perturbation sizes. We train four models in this manner. Algorithm 1 shows how we train the FGSM experts.
Input: Benign inputs: $X$, labels: $Y$, and number of epochs: $E$.
model $\leftarrow$ Instantiate(ResNet-18);
for $e\in\{0,~1,~\dots,~E-1\}$ do
 for $(x,~y)\in(X,~Y)$ do
    $\epsilon\leftarrow\textbf{Random}\{0.01,~0.02,~\dots,~0.09\}$;
    $x_{\text{adv}}\leftarrow x+\epsilon\cdot\text{sign}(\nabla_{x}L(\theta,~x,~y))$;
    $y_{p}\leftarrow\text{model.forward}(x_{\text{adv}})$;
      loss $\leftarrow\ell(y_{p},~y)$;
      model.backward(loss);
return model.
Algorithm 1 Training FGSM Experts.
We take the same strategy to train the PGD experts, with a minor difference. This time we alter the iterations number, and randomly choose a different iteration number within in each batch. With the same approach we train four PGD experts. Algorithm 2 shows how we train the PGD experts.
Input: Benign images: $X$, labels: $Y$, perturbation bound: $\epsilon$, step size: $\alpha$, and number of epochs: $E$.
model $\leftarrow$ Instantiate(ResNet-18);
for $e\in\{0,~1,~\dots,~E-1\}$ do
 for $(x,~y)\in(X,~Y)$ do
      iterations $\leftarrow\textbf{Random}\{10,~20,~30,~40,~50\}$;
    $x_{\text{adv}}\leftarrow\text{PGD}(x,~\epsilon,~\alpha,~\text{iterations})$;
    $y_{p}\leftarrow\text{model.forward}(x_{\text{adv}})$;
      loss $\leftarrow\ell(y_{p},~y)$;
      model.backward(loss);
return model.
Algorithm 2 Training PGD Experts.
Now we have a benign expert, which is trained on benign data, four FGSM experts, and four PGD experts. To train the end-to-end framework, we do not freeze the parameters of the experts. During each batch we generate FGSM and PGD version of our input, with random perturbation size and iteration number, respectively. The batch size during this stage of training is three times larger than the batch size for training the experts, which includes benign input, FGSM inputs, and PGD inputs that together can enhance clean accuracy and robustness of the network. Algorithm 3 shows the end-to-end training of our MoE network.
Input: Benign images: $X$, labels: $Y$, perturbation bound: $\epsilon$, step size: $\alpha$, number of epochs: $E$, and Experts: experts.
model $\leftarrow$ Instantiate(mixture-of-expert);
model.add(experts);
for $e\in\{0,~1,~\dots,~E-1\}$ do
 for $(x,y)\in(X,Y)$ do
    $\epsilon_{\text{FGSM}}\leftarrow\textbf{Random}\{0.01,0.02,\dots,0.09\}$;
    $x_{\text{FGSM}}\leftarrow x+\epsilon_{\text{FGSM}}\cdot\text{sign}(\nabla_{x}L(\theta,x,y))$;
      iterations $\leftarrow\textbf{Random}\{10,20,30,40,50\}$;
    $x_{\text{PGD}}\leftarrow\text{PGD}(x,\epsilon,\alpha,\text{iterations})$;
    $x_{\text{mixed}}=\text{Concatenate}(x,x_{\text{FGSM}},x_{\text{PGD}})$;
    $y_{p}\leftarrow\text{model.forward}(x_{\text{mixed}})$;
    $y_{\text{mixed}}=\text{Concatenate}(y,y,y)$;
      loss $\leftarrow\ell(y_{p},y_{\text{mixed}})$;
      model.backward(loss);
return model.
Algorithm 3 End-to-End Training of Mixture of Experts.
After all these training stages, the model is ready for testing and deployment. During both phases, DWF uses fractional schema for the experts, which enables them to contribute in the final decision. Final decision is made as follows:
$y=\sum_{i=1}^{n}w_{i}(x)\cdot f_{i}(x)$ (1)
where $n$ is the number of experts, $w_{i}(x)$ is the gating weights, and $f_{i}(x)$ stands for the $i_{th}$ expert.
5 Experiments
In this section, we firstly describe the hardware type and software libraries used in our experiments. Then, we present the dataset, target models, and adversarial attacks considered. Next, we provide details on the normalization procedure employed during both training and test phases. We also explain the learning settings used in our proposed defense system. Finally, we introduce the evaluation metrics and demonstrate the attained results.
Table 2: Accuracy and robustness of our proposed defense system vs. other models in terms of standard accuracy (SA) and robust accuracy (RA).
Model Metric SA RA - FGSM RA - PGD
0.01 0.02 0.03 0.04 0.05 0.06 0.07 0.08 0.09 10 20 30 40 50
VGG-19 93.18 52.08 50.92 48.52 41.53 29.22 20.22 16.06 13.70 12.00 42.66 41.91 41.72 41.80 41.80
ResNet-18 92.35 52.32 48.19 40.97 30.28 21.38 15.61 12.72 11.61 11.00 37.72 37.28 37.10 37.02 36.95
ResNet-50 95.29 52.68 51.54 46.01 35.04 25.46 19.26 16.25 13.87 12.14 46.20 45.68 45.53 45.44 45.45
DenseNet-169 95.10 55.32 54.57 48.63 36.38 25.21 18.57 15.00 12.88 11.90 47.55 46.99 46.81 46.72 46.81
GoogLeNet 95.01 53.83 51.53 46.96 33.59 22.70 16.61 13.83 11.69 10.91 39.60 38.59 38.53 38.48 38.49
Inception-V3 94.78 58.00 56.77 52.05 41.99 29.99 21.89 16.53 13.61 11.97 47.21 45.51 45.01 44.68 44.80
Xception 93.48 51.41 50.20 47.69 37.72 26.35 18.45 14.21 12.32 10.96 41.36 40.58 40.44 40.30 40.30
MobileNet-V2 94.20 55.05 53.79 49.31 36.67 24.70 18.31 15.38 13.39 12.19 45.77 45.20 45.14 45.17 45.08
SoE 81.69 70.23 66.22 60.27 52.65 44.62 38.26 32.96 28.74 25.74 62.27 62.34 62.31 62.31 62.31
ADVMoE 88.27 79.39 62.09 40.88 23.90 16.53 15.47 15.98 16.07 16.40 47.69 44.95 44.36 44.11 44.01
DWF (Ours) 91.08 82.83 79.66 74.76 69.01 63.60 58.40 52.97 47.34 41.24 65.51 64.7 64.52 64.42 64.37
Hardware and Implementation
All experiments were conducted on a server equipped with NVIDIA H100 GPU, offering the adequate computational power for MoE training. We implemented all the code in PyTorch with torch 2.7.1 and torchvision 0.22.0, and CUDA 12.8.
Dataset
To conduct our experiments, we use CIFAR-10 [krizhevsky2009learning] dataset, which is an RGB-colored dataset, including 10 labels.
Models
Figure 5: The backbone architecture of DWF network.
For every expert we use ResNet-18 [he2016deep] as the backbone architecture, since other relevant defenses use the same backbone. Figure 5 shows the backbone details. To compare our results with plain networks we used pretrained networks that are used in [liu2017pytorchcifar, juraev2022cifar10models]. These models include VGG-19 [simonyan2014very], ResNet-50 [he2016deep], DenseNet-169 [huang2017densely], GoogLeNet, Inception-V3 [szegedy2016rethinking], Xception [chollet2017xception], and MobileNet-V2 [sandler2018mobilenetv2]. The attained results are shown in Table 2.
Attacks
In order to train our model and test its robustness, we employ FGSM and PGD. FGSM is a single step attack, which generates adversarial samples using the following formula:
$x_{\text{adv}}=x+\epsilon\cdot sign(\nabla_{x}\ell(\theta,x,y))$ (2)
where $x$ is benign input, $y$ is the label, $x_{adv}$ is the perturbed version of $x$, $\epsilon$ is perturbation size, signsign returns the sign of the input, $\theta$ denotes the model’s parameters and $\nabla_{x}\ell(\theta,x,y)$ is the gradient of loss function with respect to $x$.
PGD, on the other hand, is an iterative attack, computing the final version of adversarial samples iteratively:
$x_{adv}^{i+1}=\prod_{x+\epsilon}\left(x_{adv}^{i}+\alpha\cdot sign\left(\nabla_{x}\ell(\theta,x,y)\right)\right)$ (3)
where $\alpha$ is step size, $i$ is the iteration number, $x_{adv}^{i}$ is the perturbed sample during the $i_{th}$ iteration, and $\prod_{x+\epsilon}$ ensures each generated adversarial sample remains within $\epsilon$ perturbation bound during each iteration. The formula for $\prod_{x+\epsilon}$ is as follows:
[…]
Table 3: Attacks settings.
Attacks Settings Attacker’s Knowledge Defense’s Knowledge
FGSM $\epsilon\in\{0.01,~0.02,...,0.09\}$ white-box white-box
PGD $\epsilon=\frac{8}{255}$, $\alpha=\frac{2}{255}$, $i\in\{10,~20,~30,~40,~50\}$ white-box white-box
Table 4: Hyper parameters - learning rate (LR), batch size (BS), and weight decay (WD).
Models Epochs LR BS WD
Experts $150,100$ $0.01$ $128,128\times 3$ 5e45e-4
Gating $100$ $0.01$ $128\times 3$ 5e45e-4
End2End $100$ $0.01$ $128\times 3$ 5e45e-4
Threat Model
To evaluate our model’s resilience against white box attacks, we employ FGSM with the perturbation sizes $\epsilon\in\{0.01,~0.02,~\dots,~0.09\}$, and PGD with the step size $\alpha=\frac{2}{255}$, perturbation bound $\epsilon=\frac{8}{255}$, and iterations numbers $i\in\{10,~20,~30,~40,~50\}$. These setting were used during both training and test phases, highlighting that both threat model and defense system have white-box knowledge about each other. Table 3 shows an overview of the attack settings.
Normalization Procedure
Let $x\in[0,1]^{32\times 32\times 3}$ be an RGB image sample from CIFAR-10. The normalization procedure for each of the three channels is as follows:
$x^{\prime}_{c}(i,j)=\frac{x_{c}(i,j)-\mu_{c}}{\sigma_{c}}$ (5)
where $c\in\{R,G,B\}$ stands for each color channel, $x^{\prime}_{c}$ is the normalized version of $x_{c}$, $\mu_{c}$ is the empirical mean over channel $c$, $\sigma_{c}$ is the empirical deviation over channel $c$, and $i,j$ are row index and column index, respectively. To normalize the data we defined $\mu=(0.4914,~0.4822,~0.4465)$ and $\sigma=(0.2023,~0.1994,~0.2010)$.
Learning Settings
We trained all the experts over $150$ epochs using batch size of $128$. We used stochastic gradient descent as the optimizer with an initial learning rate of $0.01$, momentum of $0.9$, and weight decay of $5e^{-4}$. During the training session of the end to end framework, all the models are retrained as well as the gating mechanism for $100$ epochs with batch size of $128\times 3$, where the first part includes benign inputs, the second and the third include the corresponding FGSM and PGD inputs, respectively. We also used the cross-entropy loss as the loss function for training.
Additionally, to further improve our model’s stability, we employed a cosine annealing learning rate scheduler [loshchilov2016sgdr] to decay the learning rate gradually over all iterations in all training sessions. Table 4 shows an overview of the hyper parameters used for training our proposed approach.
ADVMoE used TRADES [zhang2019theoretically] as the adversarial training objective by adopting PGD with attack strength of $\epsilon=\frac{8}{255}$ during two iterations. Additionally, it was trained during 200 epochs with learning rate of 0.1, momentum of 0.9, weight decay of $5e^{-4}$, and batch size of 128. SoE was adversarially trained during 200 epochs with learning rate of 0.1 and batch size of 128. All the experiments were done using the same backbone (ResNet-18) for DWF, ADVMoE, and SoE to ensure a fair comparison.
Metrics
To compare our model with other models, we use two metrics called standard accuracy (SA) and robust accuracy (RA). The first metric measures the model’s accuracy on the benign data, while the latter one measures the models accuracy on the adversarial data.
Figure 6: Robustness of our proposed model compared with the other competitors against FGSM w.r.t. different perturbation sizes.
Figure 7: Robustness of our proposed model vs. other models against PGD w.r.t. different iteration numbers.
Results
Table 2 illustrates the accuracy and robustness of our proposed model in comparison with other models. Figures 6 and 7 present a visual summary of the results reported in Table 2 and provide clearer comparison between our proposed model and other competitors in terms of robust accuracy under different settings for FGSM and PGD attacks. The attained results by our model are averaged over multiple runs, ensuring a fair comparison. Figures 8 and 9 show the statistics over these runs. The results over these runs vary slightly due to the randomness during training sessions. This randomness occurs due to four reasons: i) training data shuffling, ii) random weight initialization of the models before training, iii) random FGSM perturbation size of adversarial sample generation for each batch in every epoch, and iv) random PGD iteration of adversarial sample generation for each batch in every epoch.
Moreover, Figures 8 and 9 show that our proposed approach suffers from accuracy degradation, when faces stronger attacks, which is very common among all other competitors. We can see that our proposed model outperforms all other competitors in terms of robustness under all attacks. Additionally, DWF consistently outperforms existing defenses such as ADVMoE and SoE in terms of clean accuracy. Notably, it maintains a clean accuracy above $90\%$ ($91.08\pm 0.39$), highlighting its effectiveness not only against adversarial attacks but also in non-adversarial environments. Another insight illustrates that as we increase perturbation size or iterations in FGSM and PGD, the robustness of all other models drops, which shows the lack of adaptability of the models against attacks with more extreme specifications.
Figure 8: Box plots showing the robustness of our proposed approach against PGD w.r.t. different iteration numbers over multiple runs.
Figure 9: Box plots showing the robustness of our proposed approach against FGSM w.r.t. different perturbation sizes over multiple runs.
6 Conclusion and Future Works
In this paper, we proposed a novel defense system combining the ensemble architecture of MoE with adversarial training to robustify our model against PGD and FGSM attacks. Our method outperforms other competitors including undefended networks and state-of-the-art defense systems. Moreover, our proposed method, unlike the other defense systems, does not extremely sacrifice clean accuracy, underscoring its success in both adversarial and non-adversarial environments.
To reduce training time, efficient adversarial training method such as FFAT can be used. However, efficient adversarial training reduces the robustness of the model. We also trained and tested our model against two popular attacks, FGSM and PGD. To further expand our method one can train and evaluate the model on a wider range of adversarial attacks.